# Towards AI-Driven Policing: Interdisciplinary Knowledge Discovery from Police Body-Worn Camera Footage

### v2

@@ IV-C Summary Tagging with LLMs
The integration of WhisperAI, SepReformer, and Llama 3.3 with a structured data storage system not only supports current analysis but also lays the groundwork for future evaluations, refinements, and replicable studies within the OpenBWC framework.
@@ V Limitations
    Noise Distortion. BWC footage often contains significant background noise, including sirens, vehicle sounds, and overlapping speech. Although noise reduction methods made audio clearer, sometimes too much filtering removed important details like how a person sounded or subtle cues in their voice. While text transcriptions capture the words spoken, vocal details like tone and intonation carry important meaning, emotions, and intentions—details that are often lost with too much audio filtering. Finding the right balance between reducing noise and keeping these vocal details is essential for accurately understanding noisy police-civilian interactions. Additionally, the officer’s voice is often captured with greater clarity than that of civilians, likely due to the proximity of the BWC’s microphone and its orientation toward the officer’s speech. This asymmetry in audio capture can reduce the transcription quality for civilian speech in dynamic or noisy settings.
@@
    Speaker Overlap. Overlapping speech in BWC recordings complicated transcription accuracy. Although the SepReformer model effectively separated speaker audio when only two speakers were present, it struggled in chaotic scenarios characterized by rapid interruptions or when three or more speakers overlapped. This issue impacts transcription quality during precisely the types of incidents that police managers prioritize for review—high-stakes, complex interactions. This finding aligns with previous research, such as Willits and Makin (2017) [12], who reported that most of the use-of-force cases recorded by BWCs lacked sufficient audio and video clarity for human coders to reliably interpret events. Improving audio separation techniques for multi-speaker situations represents an important area for future work, given their significance in understanding police-civilian interactions.
@@ VI Future Work
Future work will expand the multi-modal approach to include voice stress detection, sentiment analysis, and object detection to identify emotionally charged segments for more targeted manual review in high-stakes scenarios. Improved diarization models may also enhance speaker identification accuracy in complex interactions. We also plan to develop cross-referencing capabilities that allow for linking multiple incidents, enhancing pattern detection, and enabling comparison across recordings of the same event captured from different BWCs. This will help address challenges posed by concatenated video releases and support a more accurate, multi-perspective understanding of dynamic police-civilian interactions. Collaboration with police officers will further support the interpretation of algorithm-clustered incidents by grounding findings in field expertise.
@@ VII Conclusion
We introduced a framework that demonstrates the potential of combining AI-driven audio, text, and image processing with advanced AI techniques for knowledge discovery from BWC footage. Analyzing repositories of BWC footage does not require access to proprietary software; instead, the tools can be open source, and the resulting algorithms, built to analyze publicly owned data, can also remain public. This promotes transparency, reproducibility, and community engagement in future research. Building on the current analysis of 1,225 videos, we identified key gaps—particularly in reliably assigning speaker identity during overlapping speech and maintaining transcription accuracy in high-noise, high-interruption scenarios. These limitations introduce uncertainty in the transcribed content. Addressing these gaps will be central to improving confidence in the outputs.
@@ Acknowledgment
This research is supported by a grant from the U.S. Department of Justice (DOJ) through its Office of Justice Programs (OJP) and Bureau of Justice Assistance (BJA), awarded to the RPD.

### v3

@@ IV-B Transcription Quality Assessment
     To support interpretability and thematic review, WhisperAI, SepReformer, and Llama 3.3 work alongside a structured data storage system. While these outputs aren’t a replacement for human judgment, they help with initial analysis and review. This integration supports current analysis and lays the groundwork for future evaluations and replicable studies within the OpenBWC framework.
@@ V Limitations
    Noise Distortion. BWC footage often contains significant background noise, including sirens, vehicle sounds, and overlapping speech. Although noise reduction methods made audio clearer, sometimes too much filtering removed important details, such as how a person sounded or subtle cues in their voice. While text transcriptions capture spoken words, vocal details such as tone and intonation convey important meaning, emotions, and intentions—details that are often lost with excessive audio filtering. Finding the right balance between reducing noise and keeping these vocal details is essential for accurately understanding noisy police-civilian interactions. Additionally, the officer’s voice is often captured with greater clarity than that of civilians, likely due to the proximity of the BWC’s microphone and its orientation toward the officer’s speech. This asymmetry in audio capture can reduce the transcription quality for civilian speech in dynamic or noisy settings.
@@
    Speaker Overlap. Overlapping speech in BWC recordings complicated transcription accuracy. Although the SepReformer model effectively separated speaker audio when only two speakers were present, it struggled in chaotic scenarios characterized by rapid interruptions or when three or more speakers overlapped. This issue impacts transcription quality during precisely the incidents police managers prioritize for review—high-stakes, complex interactions. This finding aligns with previous research, such as Willits and Makin (2017) [13], who reported that most use-of-force cases recorded by BWCs lacked sufficient audio and video clarity for human coders to reliably interpret events. Improving audio separation techniques for multi-speaker situations represents an important area for future work, given their significance in understanding police-civilian interactions.
@@ VI Future Work
Future work will use human-verified transcripts as the baseline for evaluating the system, rather than relying on automatic source separation, to improve transcription accuracy. The multimodal analysis will be refined to compare summaries of the same incident across multiple BWC videos, studying how consistent the interpretations are. We will include human-coded data to compare with automated outputs and to make the system more reliable and fair.
     Additionally, future work will expand the multi-modal approach to include voice stress detection, sentiment analysis, and object detection to identify segments with emotional content and scene recognition to analyze the visual context, helping to capture a broader range of factors influencing police-civilian interactions.
     Collaboration with police officers will further support the interpretation of algorithm-clustered incidents by grounding findings in field expertise.
     These initiatives will help strengthen the alignment between algorithmic assessments and real-world policing challenges, supporting evidence-based improvements in police training and practice.
@@ VII Conclusion
We introduced a framework demonstrating the potential of combining AI-driven audio, text, and image processing with advanced AI techniques for knowledge discovery from BWC footage. Analyzing repositories of BWC footage does not require access to proprietary software; instead, the tools can be open source, and the resulting algorithms, built to analyze publicly owned data, can also remain public. We identified key gaps in reliably assigning speaker identity during overlapping speech and maintaining transcription accuracy in high-noise, high-interruption scenarios, which introduce uncertainty in the transcribed content. Preliminary results support the value of the OpenBWC framework for interdisciplinary research, offering findings that can inform both criminological studies and practical applications in training, accountability, and policy development.
@@ Acknowledgment
This research is supported by a grant from the U.S. Department of Justice (DOJ) through its Office of Justice Programs (OJP) and Bureau of Justice Assistance (BJA), awarded to the City of Rochester.